OpenStudy (anonymous):

find the integral value of k that makes 4x^2+4x-k a perfect square trinomial, and express the result in factored form.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so this uses the discriminant. For a perfect square the discriminant is equal to zero. so you would substitute a = 4, b = 4 and k then solve for k. does that make sense..?

OpenStudy (anonymous):

would it be 4x^2+4x? if not then it doesn't make sense to me

OpenStudy (campbell_st):

ok... do you know what the discriminant is? its used to determine the type of roots that a quadratic will have..

OpenStudy (anonymous):

so that will be what0?

OpenStudy (anonymous):

if we square 4 we get 2

OpenStudy (anonymous):

i mean root will be 16

OpenStudy (campbell_st):

well start with the basics. the discriminant comes for the general quadratic formula and is \[\Delta = b^2 - 4ac\] you need a perfect square so \[\Delta = 0..... or...... 0 = b^2 - 4ac\] in your problem you have a = 4, b = 4 and c = -k you need to substitute and solve for k.

OpenStudy (anonymous):

0=b^2-16-k so far

OpenStudy (anonymous):

k=b^2-16

OpenStudy (campbell_st):

not quite its \[0 = 4^2 - 4 \times 4 \times (-k)....... or..... 0 = 16 + 16k\]

OpenStudy (anonymous):

i get k=-1

OpenStudy (anonymous):

this correct?

OpenStudy (campbell_st):

thats what I get... which doesn't make a lot of sense as the quadratic isn't a perfect square with k = -1

OpenStudy (anonymous):

oalright thanks for the help

OpenStudy (campbell_st):

glad to help...